# Cosmology using numerical relativity

### v1

@@ Abstract
This review is an up-to-date account of the use of numerical relativity to study dynamical, strong-gravity environments in a cosmological context. First, we provide a gentle introduction into the use of numerical relativity in solving cosmological spacetimes, aimed at both cosmologists and numerical relativists. Second, we survey the present body of work, focusing on general relativistic simulations without approximations, organised according to the cosmological history – from cosmogenesis, through the early hot Big Bang, to the late-time evolution of universe. In both cases, we discuss the present state-of-the-art, and suggest directions in which future work can be fruitfully pursued.
@@ 1 Introduction
$ds^{2}=-dt^{2}+a^{2}(t)(dx^{2}+dy^{2}+dz^{2})\,.$ (1)
Where such a description holds, a preferred frame exists, and we can talk meaningfully about cosmological time and space, as defined by some observers aligned with the symmetry. The Hubble parameter $H(t)=\dot{a}/a$, describing the curvature of the metric in the time direction, sets the length scale of the problem at some given time $t$ (as $L\sim H^{-1}$). The symmetry in the metric relies on assumptions about both the underlying curvature and topology of the spacetime, and the distribution of matter. In particular, present day observations have shown that matter in the universe must be roughly spatially homogeneously distributed, with no $O(1)$ overdensities on Hubble scales. Since, at the largest scales at least, this description does appear to correspond well to the current observable universe (see e.g. [1]), most cosmological simulations can be accurately approximated by splitting the background evolution of the scale factor $a(t)$ from the dynamics of the matter fields. For horizon-scale modes where the density is low and the fields are linear, a perturbative approach can be used, which is the usual approach for inflationary spacetimes (see e.g. [2]). For low density but highly non-linear field dynamics at the Hubble scale, one can numerically evolve the matter equations of motion on an expanding RW background, using the spatial average of the matter density to drive the evolution of the scale factor for consistency. If the perturbations are large, but sub-horizon and non-relativistic (i.e. local backreaction matters), then the dynamics can be approximated using Newtonian gravity on an expanding RW background, a strategy that is the leading approach in large-scale N-body simulations of late-time cosmological evolution (see e.g. [3, 4]).
While numerical relativity (NR) is a standard tool in the computation of accurate gravitational waveforms in compact object astrophysics (i.e. black holes (BHs) and neutron stars (NSs)), its application to cosmology is not yet widespread. Recent works that implement improvements in gauge choices, initial conditions and stability have nevertheless been pushing forward our understanding of the universe, building on a number of pioneering early explorations in the field. Cosmologists are now beginning to harness the considerable power of NR to explore strong gravity problems that are beyond the reach of standard perturbative treatments.
@@
In this review, we focus on fully general relativistic simulations without approximations (i.e. not assuming a certain background or perturbative description). We provide useful background in Sec. 2. Recognising that most cosmologists are not familiar with the methods of numerical relativists, and vice versa, we start by providing two reviews of the topic from different angles – first we explain cosmology to numerical relativists, and then we explain NR to cosmologists. We then review common technical and conceptual challenges of using NR in cosmological applications, which are themselves open topics of research. The resolution of some of these challenges may lead to new techniques that can be used in other applications of NR, and also to deeper insights about the nature of the cosmological spacetime in which we exist.
@@ 2.1 Cosmology for numerical relativists
We provide here only a brief description of cosmology in the context of NR. A good starting point for those new to cosmology are textbooks by Baumann [2], Weinberg [14] and Mukhanov [15].
@@
known as the Hubble constant or, less misleadingly, the Hubble parameter since it is a function of cosmological time. Currently, there exists an observational tension in measurements of its exact present day value $H_{0}$ (read as H-nought) – measurement of $H_{0}$ from observations of the Cosmic Microwave Background (CMB) and the clustering of galaxies suggest that $H_{0}\approx 68~{}\mathrm{km}\mathrm{s}^{-1}~{}\mathrm{Mpc}^{-1}$ [1, 16], while observations of astrophysical standard candles suggest a value of $H_{0}$ between $H_{0}\approx 73~{}\mathrm{km}\mathrm{s}^{-1}~{}\mathrm{Mpc}^{-1}$ and $H_{0}\approx 70~{}\mathrm{km}\mathrm{s}^{-1}~{}\mathrm{Mpc}^{-1}$ [17, 18]. Importantly, cosmological measurements are not direct measurements of the present day value but inferred values from measurements at a different epoch, assuming a certain subsequent cosmological evolution. So the tension potentially provides evidence that the standard $\Lambda$CDM model (dark energy $\Lambda$ plus cold dark matter) is incomplete.
@@
with $U_{\mu}=(1,0,0,0)$ the local 4-velocity of a co-moving fluid parcel. The first equation of Eqn. (3) can be recognised as the Hamiltonian constraint, while the second equation describes the evolution of $H=\dot{a}/a$, which is equivalent to the evolution equation for the trace of the extrinsic curvature $K$ in the ADM formalism described in the next section.
@@
$p=w(t)\rho~{},$ (6)
where $w(t)$ is known as the equation of state parameter, or simply double-u. This parameterisation, whilst only an approximation for most actual fluids, allows cosmologists to study the impact of different components of the matter content on the cosmological evolution, i.e. $w=0$ for pressureless matter, $w=1/3$ for radiation and $w=-1$ for the cosmological constant. Where multiple matter types are present they are usually assumed independent and must be evolved separately, with their densities and pressures added to obtain the sources in Eqn. (3).
@@
and $h_{\mu\nu}\ll g_{\mu\nu}^{(0)}$. Finally, the 3D spatial metric $\gamma_{ij}$ is maximally symmetric. Whilst one can include a spatially constant curvature, for simplicity we will consider only foliations that are spatially flat $\gamma_{ij}=\delta_{ij}$ for the remainder of this section.
In the last two decades or so, cosmological perturbation theory (the study of the structure and evolution of cosmological perturbations) has yielded a remarkable amount of information about the universe, heralding the era of Precision Cosmology (see Turner [21] for a historical overview). The study is statistical in nature – e.g. it is not important where individual galaxies are, but it is important how the separations of galaxies are statistically correlated with each other. The key observables of structure are correlation functions, $\langle X_{1}X_{2}\dots\rangle$ where $X_{i}$ are cosmological observables such as density perturbations or polarisation amplitudes, and the angle brackets denote an average over spatial hyperslicings. The most important quantity is the power spectrum or two-point correlation function – for example the matter power spectrum, which is the two-point correlation of the density perturbation $\langle\delta\delta\rangle$ where $\delta=\delta\rho/\rho$. As long as the perturbations remain small, in principle correlation functions at all orders can be calculated via perturbation theory (see for example [22, 23]).
@@
$G^{(0)}_{\mu\nu}+\delta G_{\mu\nu}=8\pi G(T^{(0)}_{\mu\nu}+\delta T_{\mu\nu})~{}.$ (9)
We first solve the background evolution $G^{(0)}_{\mu\nu}=8\pi GT^{(0)}_{\mu\nu}$ to obtain the zeroth order terms for both the metric and the matter sector, and then plug those into the linearised Einstein equations $\delta G_{\mu\nu}=8\pi G\delta T_{\mu\nu}$, which describes the inhomogeneous dynamics of the spacetime. We similarly linearise the matter equations of motion, which combined with Eqn. (9), will close the system of equations consistently. For the details of this (long but straightforward) calculation, see for example [2].
We now clarify some of the nomenclature surrounding cosmological gauges, which can be confusing to non-cosmologists approaching it from an NR perspective. In GR, gauge freedom refers to the freedom to choose coordinates, such that any physical measurement should remain invariant under a coordinate transform. In general, this transformation can be non-linear, and applies to the whole spacetime metric and its slicing. In linear perturbation theory on a fixed cosmological background, this freedom is translated into the statement that any two perturbations $h_{\mu\nu}$ and $h^{\prime}_{\mu\nu}$ are physically the same as long as they are related to each other by a gauge transformation $\Delta h_{\mu\nu}$, which is a function of the background metric $g^{(0)}_{\mu\nu}$ and a gauge choice. Specifically, suppose we have a linear coordinate transform [^2] Note that while the coordinate $x^{\mu}$ is not a spacetime vector, the infinitisimal transform $\epsilon^{\mu}$ is and hence its indices can be raised and lowered as usual.
@@
All perturbations that are related by this gauge transform are physically equivalent. Thus there exist a redundancy in the description of the perturbations that one can either remove by fixing a gauge (that is, choosing a specific $\epsilon^{\mu}$), or by choosing to work with gauge-invariant variables [19, 24, 25].
@@
To summarise, in cosmological perturbation theory, the background metric $g^{(0)}_{\mu\nu}$ determines both the foliation and the coordinates on each hyperslice with normal $n^{\mu}=\partial_{t}$ – the choice is already made here [^3] One can of course foliate the RW spacetime with inhomogeneous hyperslicings where the symmetry is not made manifest, but this would require a change of coordinates which mix up $t$ and $x^{i}$.. Gauge freedom is restricted to linear transforms on the perturbation fields $h_{\mu\nu}$ and the corresponding matter stress tensor $\delta T_{\mu\nu}$, which manifests as a redundancy in the description of the perturbation variables. Contrast this to the fully non-linear case in an Arnowitt-Deser-Misner [27] (ADM) decomposition, as used in NR, where gauge-fixing refers to the choice of the lapse $\alpha$ and shift $\beta^{i}$ functions, and fixes the temporal foliation and the spatial coordinates as described in Sec. 2.2 and 2.5. The difference is illustrated in Fig. 3. What cosmologists describe as scalar, vector and tensor perturbations do not transform in the expected tensorial way with changes of the ADM gauge, and moreover cosmologists think of scalars as gauge dependent and tensors as gauge invariant. This standard interpretation is problematic when we want to compare results between perturbation theory and NR, and in fact there is no obvious mapping between the two pictures once perturbations become large relative to the background. One must reformulate the question in an observer dependent way – in terms of what particular geodesic observers moving through each spacetime would measure. This picture is non-trivial to reconstruct, and so less rigorous diagnostics like spatial averages are often used as a proxy, as will be discussed in Sec 2.6.
Figure 3: Gauge in cosmology versus NR. (Left:) For cosmologists a change in gauge implies a small perturbation to the slicing, but with the background FLRW slice about which the perturbation is defined always aligned with the symmetry of the spacetime. This can transform a constant density slice into one in which small perturbations are observed. (Right:) In the NR case, a change of gauge is a fundamentally different slicing of the spacetime, which therefore potentially results in a (non-perturbative) change in the density observed on each constant-time slice. It is important to note that when there is no underlying symmetry, the notion of a preferred slicing is lost, and thus there is no choice of gauge in which we are able to recover FLRW-like observers.
@@ 2.2 Numerical relativity for cosmologists
The basic principle of NR is to solve the coupled matter-metric system of GR as a Cauchy or initial value problem, whereby we specify initial data for the matter and the metric on some space-like hypersurface, and then evolve it forward in time (defined according to some chosen observers); that is, we specify the initial scenario, and then use the computer to find out what happens next. Such a formulation is naturally provided by the ADM decomposition of the Einstein equations, which was developed in the 1960s by Arnowitt, Deser, and Misner [27].
@@
A standard question when confronted with any formulation of PDEs is whether they are well-posed, which is the demand that a unique solution exists that depends continuously on the initial data. In the perturbative cosmology of the previous Sec. 2.1, one separates the background evolution and the perturbation evolution equations, both of which are well-posed. In fully non-linear NR, the well-posedness of the formulation is less obvious and problems with well-posedness in NR have a long history. In particular, one of the early obstacles in NR was that the ADM formulation [27] was ill-posed, rendering its direct numerical implementation unstable. Fortunately, this problem can be surmounted by reformulating the evolution equations into one of several possible (locally) well-posed forms. For a detailed introduction to well-posedness in NR, see the review article by Hilditch [33]. One popular formulation is the so-called Baumgarte-Shapiro-Shibata-Nakamura (BSSN) formalism [34, 35] based also on an earlier work with Nakamura, Oohara and Kojima [36]. An alternative formulation is the Generalised Harmonic Coordinates (GHC) formalism [37], which is well-known for its use in successfully evolving binary BH mergers for the first time by Pretorius [38], but was applied to studies of the cosmological singularity before this by Garfinkle [39]. Another alternative is the tetrad formulation developed by Uggla et al. [40], used by Garfinkle for the study of cosmological singularities [41] and recent works on bouncing cosmologies (see Sec. 3.3), which allows scaling out of the cosmological scale $H$. This helps the numerics to follow longer timescales during which $H$ is changing over a wide range of scales, as at the end of inflation or contraction towards a singularity, but necessitates the solution of an elliptic equation for the lapse at each step. There are some other early works that formulate hyperbolic equations with tetrads [42, 43], which were mostly abandoned when coordinate basis formulations achieved success with binary BH mergers, but may be worth revisiting to address some of the challenges of cosmology.
@@
In this section we will mainly discuss NR in the context of the BSSN formalism, but the principles are similar for most formulations. The idea of slicing up spacetime into space-like hypersurfaces and then evolving them is familiar to most cosmologists, where the RW metric imposes a fixed foliation and coordinates upon each hyperslice as discussed in Sec. 2.1. In NR, the choice of foliation and the coordinates imposed on it are usually dynamical – they respond to the local evolution of the matter and metric. This is a necessity for a well-posed evolution, but can also help with avoiding singularities that form and tracking changing scales over time on a finite numerical grid.
@@
together with $\gamma_{ij}$, forms the phase space of metric degrees of freedom.
@@
which must be satisfied on every hypersurface. The decomposed components of the matter stress-energy tensor (those measured by normal observers) are defined as
@@
With the degrees of freedom and their equations of motion defined, the system can in principle be discretised and solved on a computer using standard numerical integration and differentiation techniques (see e.g. the relevant chapters of [29, 30]). However, as discussed above, due to the additional complication of ensuring the equations are well-posed, the ADM equations are modified and auxilliary variables may be defined depending on the formulation, see Hilditch [33] for a review.

### v2

@@ Abstract
This review is an up-to-date account of the use of numerical relativity to study dynamical, strong-gravity environments in a cosmological context. First, we provide a gentle introduction into the use of numerical relativity in solving cosmological spacetimes, aimed at both cosmologists and numerical relativists. Second, we survey the present body of work, focusing on general relativistic simulations, organised according to the cosmological history – from cosmogenesis, through the early hot Big Bang, to the late-time evolution of universe. We discuss the present state-of-the-art, and suggest directions in which future work can be fruitfully pursued.
@@ 1 Introduction
$ds^{2}=-dt^{2}+a^{2}(t)(dx^{2}+dy^{2}+dz^{2})\,,$ (1)
assumming $c=1$. Where such a description holds, a preferred frame exists, and we can talk meaningfully about cosmological time and space, as defined by some observers aligned with the symmetry. The Hubble parameter $H(t)=\dot{a}/a$, describing the curvature of the metric in the time direction, sets the length scale of the problem at some given time $t$ (as $L\sim H^{-1}$). The symmetry in the metric relies on assumptions about both the underlying curvature and topology of the spacetime, and the distribution of matter. In particular, present-day observations [1] provide direct evidence for isotropy, which combined with the Copernican principle imply that matter in the universe must be roughly spatially homogeneously distributed, with no $O(1)$ overdensities on Hubble scales [^1] See however, [2, 3] for caveats on inferring the homogeneity of the universe via observations along the lightcone.. Since, at the largest scales at least, this description does appear to correspond well to the current observable universe, most cosmological simulations can be accurately approximated by splitting the background evolution of the scale factor $a(t)$ from the dynamics of the matter fields. For horizon-scale modes where the density is low and the fields are linear, a perturbative approach can be used, which is the usual approach for inflationary spacetimes (see e.g. [4]). For low density but highly non-linear field dynamics at the Hubble scale, one can numerically evolve the matter equations of motion on an expanding RW background, using the spatial average of the matter density to drive the evolution of the scale factor for consistency. If the perturbations are large, but sub-horizon and non-relativistic (i.e. local backreaction matters), then the dynamics can be approximated using Newtonian gravity on an expanding RW background, a strategy that is the leading approach in large-scale N-body simulations of late-time cosmological evolution (see e.g. [5, 6]).
While numerical relativity (NR) is a standard tool in the computation of accurate gravitational waveforms in compact object astrophysics (i.e. black holes (BHs) and neutron stars (NSs)), its application to cosmology is not yet widespread. Nevertheless, numerical investigations in cosmological spacetimes have a long history, and some of the earliest works, mostly focussed on 1D planar cosmologies, deserve to be highlighted as pioneering work in the field [7, 8, 9, 10, 11, 12, 13, 14]. In particular, works by Centrella, Matzner and their collaborators explored a wide range of topics including the collision of gravitational waves [15], planar shocks [16] and nucleosynthesis [17]. More recent works that implement improvements in gauge choices, initial conditions and stability have built on these foundations to push forward our understanding of the universe. Cosmologists are now beginning to harness the considerable power of NR to explore strong gravity problems that are beyond the reach of standard perturbative treatments.
@@
In this review, we focus on fully general relativistic simulations without approximations (i.e. not assuming a certain background, conformal flatness or perturbative or post-Friedmann description). This technically excludes some interesting and useful intermediate approaches between the non-linear and linear regimes that have been suggested in both NR and cosmology, e.g. [18, 19, 20, 21, 22, 23, 24], that we nevertheless briefly highlight where relevant. We provide useful background in Sec. 2. Recognising that most cosmologists are not familiar with the methods of numerical relativists, and vice versa, we start by providing two reviews of the topic from different angles – first we explain cosmology to numerical relativists, and then we explain NR to cosmologists. We then review common technical and conceptual challenges of using NR in cosmological applications, which are themselves open topics of research. The resolution of some of these challenges may lead to new techniques that can be used in other applications of NR, and also to deeper insights about the nature of the cosmological spacetime in which we exist.
@@ 2.1 Cosmology for numerical relativists
We provide here only a brief description of cosmology in the context of NR. A good starting point for those new to cosmology are textbooks by Baumann [4], Weinberg [34], Mukhanov [35], Ryden [36], Peacock [37], Ellis, Maartens & MacCallum [38] and Huterer [39].
@@
which is clearly a variable that is a function of time but is often referred to as the Hubble parameter, or when discussing its present day value the Hubble constant. Currently, there exists an observational tension in measurements of its exact present day value $H_{0}$ (read as H-nought) – measurement of $H_{0}$ from observations of the Cosmic Microwave Background (CMB) and the clustering of galaxies suggest that $H_{0}=67.97\pm 0.38~{}\mathrm{km}\mathrm{s}^{-1}~{}\mathrm{Mpc}^{-1}$ [1, 40], while observations of astrophysical standard candles suggest $H_{0}=73.30\pm 1.04~{}\mathrm{km}\mathrm{s}^{-1}~{}\mathrm{Mpc}^{-1}$ [41, 42]. Importantly, cosmological measurements are not direct measurements of the present day value but inferred values from measurements at a different epoch, assuming a certain subsequent cosmological evolution. So the tension potentially provides evidence that the standard $\Lambda$CDM model (dark energy $\Lambda$ plus cold dark matter) is incomplete.
@@
with $U_{\mu}=(1,0,0,0)$ the local 4-velocity of a co-moving fluid parcel. The first equation of Eqn. (3) can be recognised as the Hamiltonian constraint, while the second equation describes the evolution of $H=\dot{a}/a$, which is the Raychaudhuri expansion equation in the FLRW limit. As we will see in the next section, this is equivalent to the evolution equation for the trace of the extrinsic curvature $K=-3H$ in the ADM formalism.
@@
$p=w\rho~{},$ (6)
where $w$ is known as the equation of state parameter, or simply double-u [^2] This parameter can in principle also depend on time, although in this case the equations do not close unless some other ansatz is made, for example assuming a dependence on the scale factor like that currently used to parameterise an evolving dark energy $w(a)=w_{0}+w_{a}(1-a)$, which was first introduced in [43].. This parameterisation, whilst only an approximation for most actual fluids, allows cosmologists to study the impact of different components of the matter content on the cosmological evolution, i.e. $w=0$ for pressureless matter, $w=1/3$ for radiation and $w=-1$ for the cosmological constant. Where multiple matter types are present they are usually assumed independent and must be evolved separately, with their densities and pressures added to obtain the sources in Eqn. (3).
@@
and $|h_{\mu\nu}|\ll g_{\mu\nu}^{(0)}$. Finally, the 3D spatial metric $\gamma_{ij}$ is maximally symmetric. Whilst one can include a spatially constant curvature, for simplicity we will consider only foliations that are spatially flat $\gamma_{ij}=\delta_{ij}$ for the remainder of this section.
Cosmological perturbation theory (the study of the structure and evolution of cosmological perturbations) has yielded a remarkable amount of information about the universe, heralding the era of Precision Cosmology (see Turner [46] for a historical overview). The study is statistical in nature – e.g. it is not important where individual galaxies are, but it is important how the separations of galaxies are statistically correlated with each other. The key observables of structure are correlation functions, $\langle X_{1}X_{2}\dots\rangle$ where $X_{i}$ are cosmological observables such as density perturbations or polarisation amplitudes, and the angle brackets denote an average over spatial hyperslicings. The most important quantity is the power spectrum or two-point correlation function – for example the matter power spectrum, which is the two-point correlation of the density perturbation $\langle\delta\delta\rangle$ where $\delta=\delta\rho/\rho$. As long as the perturbations remain small, in principle correlation functions at all orders can be calculated via perturbation theory (see for example [47, 48]).
@@
$G^{(0)}_{\mu\nu}+\delta G_{\mu\nu}=8\pi G(T^{(0)}_{\mu\nu}+\delta T_{\mu\nu})~{},$ (9)
We first solve the background evolution $G^{(0)}_{\mu\nu}=8\pi GT^{(0)}_{\mu\nu}$ to obtain the zeroth order terms for both the metric and the matter sector, and then plug those into the linearised Einstein equations $\delta G_{\mu\nu}=8\pi G\delta T_{\mu\nu}$, which describes the inhomogeneous dynamics of the spacetime. We similarly linearise the matter equations of motion, which combined with Eqn. (9), will close the system of equations consistently. For the details of this (long but straightforward) calculation in the case of a cosmological background, see for example [4], and also the reviews of [49, 50] for other approaches, and going beyond the linear regime.
We now clarify some of the nomenclature surrounding cosmological gauges, which can be confusing to non-cosmologists approaching it from an NR perspective. In GR, gauge freedom refers to the freedom to choose coordinates, such that any physical measurement should remain invariant under a coordinate transform. In general, this transformation can be non-linear, and applies to the whole spacetime metric and its slicing. In linear perturbation theory on a fixed background, this freedom is translated into the statement that any two perturbations $h_{\mu\nu}$ and $h^{\prime}_{\mu\nu}$ are physically the same as long as they are related to each other by a gauge transformation $\Delta h_{\mu\nu}$, which is a function of the background metric $g^{(0)}_{\mu\nu}$ and a gauge choice [^4] In this discussion, we have chosen to follow the treatment of standard cosmology texts, which usually focuses on the more explicit coordinate-centric approach. There is an elegant geometrical picture to describe this equivalence, which is perhaps more familiar to the study of gravitational radiation as follows. One posits the existence of two spacetime manifolds which are related by a diffeomorphism $\phi:{\cal M}_{b}\rightarrow{\cal M}_{p}$, where ${\cal M}_{b}$ is the “background spacetime” and ${\cal M}_{p}$ is the “physical spacetime”, with different tensor fields living on them. On ${\cal M}_{b}$ we define the “background” metric $g_{\mu\nu}^{(0)}$, and while on ${\cal M}_{p}$, we define the “physical” metric $g_{\mu\nu}$. One can then treat linearized theory as the behaviour of tensor field perturbations $h_{\mu\nu}$ on ${\cal M}_{b}$, where $h_{\mu\nu}$ is the difference between the pull-back metric $\phi_{*}(g)_{\mu\nu}$ from ${\cal M}_{p}$ and the background spacetime metric, i.e. $h_{\mu\nu}=\phi_{*}(g)_{\mu\nu}-g_{\mu\nu}^{(0)}$ (notice the placement of the indices outside the bracket in the pullback of $g$). In this picture, the gauge freedom at first order is generated by the choice of vector field $\epsilon^{\mu}$ on ${\cal M}_{b}$, and hence $h_{\mu\nu}$ in different gauges are physically equivalent. For gauge transformations at higher orders, please see [51, 52]. . Specifically, suppose we have a linear coordinate transform [^5] Note that while the coordinate $x^{\mu}$ is not a spacetime vector, the infinitisimal transform $\epsilon^{\mu}$ is and hence its indices can be raised and lowered as usual.
@@
All perturbations that are related by this gauge transform are physically equivalent. Thus there exists a redundancy in the description of the perturbations that one can either remove by fixing a gauge (that is, choosing a specific $\epsilon^{\mu}$), or by choosing to work with gauge-invariant variables [44, 53, 54].
@@
To summarise, in cosmological perturbation theory, the background metric $g^{(0)}_{\mu\nu}$ determines both the foliation and the coordinates on each hyperslice with normal $n^{\mu}=\partial_{t}$ – the choice is already made here [^6] One can of course foliate the RW spacetime with inhomogeneous hyperslicings where the symmetry is not made manifest, but this would require a change of coordinates which mix up $t$ and $x^{i}$.. Gauge freedom is restricted to linear transforms on the perturbation fields $h_{\mu\nu}$ and the corresponding matter stress tensor $\delta T_{\mu\nu}$, which manifests as a redundancy in the description of the perturbation variables. Contrast this to the fully non-linear case in an Arnowitt-Deser-Misner [56] (ADM) decomposition, as used in NR, where gauge-fixing refers to the choice of the lapse $\alpha$ and shift $\beta^{i}$ functions, and fixes the temporal foliation and the spatial coordinates as described in Sec. 2.2 and 2.5. The difference is illustrated in Fig. 3.
In summary, when cosmologists describe scalar, vector and tensor perturbations, they refer to the invariant transformation properties of spatial tensors and not spacetime tensors [44]. Thus as a consequence scalar perturbations are gauge dependent and tensors are gauge invariant under linear spacetime coordinate transformations.
This standard interpretation is problematic when we want to compare results between perturbation theory and NR, and in fact there is no obvious mapping between the two pictures once perturbations become large relative to the background. One must reformulate the question in an observer dependent way – in terms of what particular geodesic observers moving through each spacetime would measure. This picture is non-trivial to reconstruct, and so less rigorous diagnostics like spatial averages are often used as a proxy, as will be discussed in Sec 2.6.
Figure 3: Gauge in cosmology versus NR. (Left:) For cosmologists a change in gauge implies a small perturbation to the slicing, but with the background FLRW slice about which the perturbation is defined always aligned with the symmetry of the spacetime. This can transform a constant density slice into one in which small perturbations are observed. (Right:) In the NR case, a change of gauge is a fundamentally different slicing of the spacetime (conventionally parameterised by a choice of the lapse and shift), which therefore potentially results in a (non-perturbative) change in the density observed on each constant-time slice. It is important to note that when there is no underlying symmetry, the notion of a preferred slicing is lost, and thus there is no choice of gauge in which we are able to recover FLRW-like observers.
@@ 2.2 Numerical relativity for cosmologists
The basic principle of NR is to solve the coupled matter-metric system of GR as a Cauchy or initial value problem, whereby we specify initial data for the matter and the metric on some space-like hypersurface, and then evolve it forward in time (defined according to some chosen observers); that is, we specify the initial scenario, and then use the computer to find out what happens next. Such a formulation is naturally provided by the ADM decomposition of the Einstein equations, which was developed in the 1960s by Arnowitt, Deser, and Misner [56] as a Hamiltonian formulation of GR.
@@
A standard question when confronted with any formulation of PDEs is whether they are well-posed, which is the demand that a unique solution exists that depends continuously on the initial data. In the perturbative cosmology of the previous Sec. 2.1, one separates the background evolution and the perturbation evolution equations, both of which are well-posed. In fully non-linear NR, the well-posedness of the formulation is less obvious and problems with well-posedness in NR have a long history. In particular, one of the early obstacles in NR was that the need for a well posed system of equations was not sufficiently appreciated, and since the ADM formulation [56] was ill-posed, its direct numerical implementation was unstable. This problem can be surmounted by reformulating the evolution equations into one of several possible (locally) well-posed forms. For a detailed introduction to well-posedness in NR, see the review article by Hilditch [62]. One popular reformulation of the ADM system is the so-called Baumgarte-Shapiro-Shibata-Nakamura (BSSN) formalism [63, 64] based also on an earlier work with Nakamura, Oohara and Kojima [65]. An alternative formulation is Generalised Harmonic Coordinates (GHC), which was introduced by Friedrich [66] in 1985 and based on the harmonic coordinates that had been shown to be well posed by Choquet-Bruhat already in 1952 [67]. GHC is well-known for its use in successfully evolving binary BH mergers for the first time by Pretorius [68], but was applied to studies of the cosmological singularity before this by Garfinkle [69]. Another alternative is the tetrad formulation developed by Uggla et al. [70], used by Garfinkle for the study of cosmological singularities [71] and recent works on bouncing cosmologies (see Sec. 3.3), which allows scaling out of the cosmological scale $H$. This helps the numerics to follow longer timescales during which $H$ is changing over a wide range of scales, as at the end of inflation or contraction towards a singularity, but necessitates the solution of an elliptic equation for the lapse at each step. There are some other early works that formulate hyperbolic equations with tetrads [72, 73], which were mostly abandoned when coordinate basis formulations achieved success with binary BH mergers, but may be worth revisiting to address some of the challenges of cosmology.
@@
In this section we will mainly discuss NR in the context of the ADM formalism, focussing on the broad principles, which are similar for most of the more well posed formulations. However, anyone hoping to use NR codes in practise will need to familiarise themselves with one of the above formulations – to keep this review focussed we leave a discussion of these to the NR textbooks mentioned above. The idea of slicing up spacetime into space-like hypersurfaces and then evolving them is familiar to most cosmologists, where the RW metric imposes a fixed foliation and coordinates upon each hyperslice as discussed in Sec. 2.1. In NR, the choice of foliation and the coordinates imposed on it are usually dynamical – they respond to the local evolution of the matter and metric. This is a necessity for a well-posed evolution, but can also help with avoiding singularities that form and tracking changing scales over time on a finite numerical grid.
@@
(where $D_{i}\equiv\gamma_{i}^{\mu}\nabla_{\mu}$) together with $\gamma_{ij}$, forms the phase space of metric degrees of freedom.
@@
which must be satisfied on every hypersurface. Here ${}^{(3)}R_{ij}$ is the Ricci tensor associated to the 3-metric, and the decomposed components of the matter stress-energy tensor (those measured by normal observers) are defined as
@@
With the degrees of freedom and their equations of motion defined, the system can in principle be discretised and solved on a computer using standard numerical integration and differentiation techniques. However, as discussed above, due to the additional complication of ensuring the equations are well-posed, the ADM equations are not suitable for evolving (see Hilditch [62] for a review). These are modified, and auxiliary variables may be defined depending on the formulation (see e.g. the relevant chapters of [58, 59]).